There are two identical, positively charged conducting spheres fixed in space. The spheres are 31.6 cm apart (center to center) and repel each other with an electrostatic force of F1 = 0.0795 N. Then, a thin conducting wire connects the spheres, redistributing the charge on each sphere. When the wire is removed the spheres still repel but with a force of F2 = 0.100 N. Using this information, find the initial charge on each sphere, q1 and q2 if initially q1. Please answer with numerical answers and explanation. Thank you.
Answer #1

After the spheres are connected, the charges redistribute so there is the same (positive) charges, Q, on each sphere

F= kq1q2/r^2

0.1 = 9*10^9 * Q * Q /(0.316^2)

Q = sqrt(1.11*10^(-12)) = 1.053*10^(-6) C

now total charge on both sphere = 2*Q = 2.106*10^(-6) C

so q1 + q2 = 2.106*10^(-6) C

now before seperation

0.0795 = 9*10^9 * q1 * q2 /(0.316^2)

q1 * q2 = 8.82*10^(-13) C^2

q2 = 8.82*10^(-13) / q1

q1 + 8.82*10^(-13) / q1 = 2.106*10^(-6)

q1^2 - 2.106*10^(-6)*q1 + 0.882*10^(-12) = 0

q1 = 1.53*10^(-6) C

q2 = 0.576*10^(-6) C